# Temperature Dependent Spin-Phonon Coupling of Boron-Vacancy Centers in Hexagonal Boron Nitride

### v1

@@ Abstract
Dynamic nuclear polarization at low temperature— With the temperature-dependent $\mathrm{V}_{\mathrm{B}}^{-}$ spin properties in hand, we now turn to investigate the dynamical polarization of the three nearest-neighbor ${}^{15}\mathrm{N}$ nuclear spins at low temperature. Nuclear spins feature exceptional isolation from noisy environments, making them ideal candidates for quantum storage applications bradley2019ten ; metsch2019initialization . At room temperature, several prior works have demonstrated the polarization of proximate 14N or 15N with an upper limit of $\sim 30~{}\%$; however how to further improve the polarization fidelity remains an open question  gong2024isotope ; gao2022nuclear ; ru2023robust . Low temperature offers two potential advantages: $\sim 100$x longer $\mathrm{V}_{\mathrm{B}}^{-}$ electronic spin lifetime $T_{1}$ and stronger hyperfine interaction strength.
@@
To enable resonant spin exchange between electronic and nuclear spins (known as electronic spin level anti-crossing, esLAC) gong2024isotope ; gao2022nuclear ; mathur2022excited , we apply an external magnetic field $B_{z}\approx 760~{}$G under which the energy levels corresponding to $\mathrm{V}_{\mathrm{B}}^{-}$ excited state $|m_{s}=0\rangle$ and $|m_{s}=-1\rangle$ are nearly degenerate (see Supplementary Materials). We fit the measured hyperfine resonances between $\mathrm{V}_{\mathrm{B}}^{-}$ $|m_{s}=0\rangle\leftrightarrow|m_{s}=+1\rangle$ in the ODMR spectra as a function of the laser power to extract the nuclear spin polarization fidelity (Fig. 4a). We find that the nuclear spin polarization at both $10$ K and 300300300~{}300K saturate around $30\%$ with similar laser powers ($\sim 10~{}$mW) (Fig. 4b), suggesting that the $\mathrm{V}_{\mathrm{B}}^{-}$ spin lifetime $T_{1}$ and hyperfine interaction strength are not the limiting sources for nuclear polarization.
@@
Figure 4: Dynamic nuclear polarization at low temperature (a) Near esLAC level ODMR Spectra of the $|m_{s}=0\rangle\leftrightarrow|m_{s}=-1\rangle$ transition at 10 K and 300 K. The spectra both exhibit similar asymmetry toward the left peaks, indicating a polarization of nuclear spins (see Supplementary Materials) (b) Polarization level of the three nearest ${}^{15}\mathrm{N}$ nuclear spins near esLAC level as a function of the normalized laser power under room/cryogenic temperature. Inset: schematic representation of the polarization process. The strong spin-conserving optical polarization (green arrow) continuously pumps state from $|m_{s}=-1\rangle$ to $|m_{s}=0\rangle$, and the two hybridisation processes (red arrow and dashed blue arrow) differ in strength, resulting in the polarization of $|m_{I}=\>\uparrow\rangle$ state (see Supplementary Materials).
@@
The spin polarization may be limited from the spin non-conserving terms in the hyperfine interaction (see Supplementary Materials), which are temperature insensitive. In addition, the observed nuclear spin polarization of $\mathrm{V}_{\mathrm{B}}^{-}$ persists down to 101010~{}10K, in sharp contrast to the case of NV centers where the dynamic nuclear polarization process is substantially suppressed at $T\lesssim 50~{}$K, due to the lack of thermal averaging between NV excited states originating from different orbitals fischer2013optical ; block2021optically . For $\mathrm{V}_{\mathrm{B}}^{-}$ , the lifting of the double degeneracy of excited states leads to a sufficiently large energy gap between the states, which renders the system insensitive to dynamic Jahn–Teller distortions mathur2022excited .
@@
Conclusion and Outlook— In summary, our research offers a comprehensive investigation of the temperature-dependent spin-phonon interaction of $\mathrm{V}_{\mathrm{B}}^{-}$ centers in isotopically purified $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$. Our results will be of critical importance in designing experiments to make use of $\mathrm{V}_{\mathrm{B}}^{-}$ centers as nanoscale thermometers and phonon sensors. Moreover, although we focus on $\mathrm{V}_{\mathrm{B}}^{-}$, the methodologies developed here can be readily applied to a broad family of spin defects in two-dimensional materials stern2023quantum ; guo2023coherent ; li2022carbon ; lee2022spin .
@@
Acknowledgements: We gratefully acknowledge Vincent Jacques, Tongcang Li, Shankar Mukherji, Erdong Song, Xi Wang and Norman Yao for helpful discussions. This work is supported by NSF ExpandQISE 2328837 and the Center for Quantum Leaps of Washington University. B. Huang, Y. Jin and G. Galli acknowledge the support of Air Force Office of Scientific Research (AFOSR) under Contract No. FA9550-22-1-0370, the Next Generation Quantum Science and Engineering (QNEXT) hub supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and resources of the University of Chicago Research Computing Center. Support for hBN crystal growth (E. Janzen and J. H. Edgar) was provided by the Office of Naval Research, award no. N00014-22-1-2582.

### v2

@@ Abstract
Dynamic nuclear polarization at low temperature— With the temperature-dependent $\mathrm{V}_{\mathrm{B}}^{-}$ spin properties in hand, we now turn to investigate the dynamical polarization of the three nearest-neighbor ${}^{15}\mathrm{N}$ nuclear spins at low temperature. Nuclear spins feature exceptional isolation from noisy environments, making them ideal candidates for quantum storage applications [73, 74, 75]. At room temperature, several prior works have demonstrated the polarization of proximate 14N or 15N with an upper limit of $\sim 30~{}\%$; however how to further improve the polarization fidelity remains an open question [49, 76, 77]. Low temperature offers two potential advantages: $\sim 100$x longer $\mathrm{V}_{\mathrm{B}}^{-}$ electronic spin lifetime $T_{1}$ and stronger hyperfine interaction strength.
@@
To enable resonant spin exchange between electronic and nuclear spins (known as electronic spin level anti-crossing, esLAC) [49, 76, 78], we apply an external magnetic field $B_{z}\approx 760~{}$G under which the energy levels corresponding to $\mathrm{V}_{\mathrm{B}}^{-}$ excited states $|m_{s}=0\rangle$ and $|m_{s}=-1\rangle$ are nearly degenerate (see Supplemental Material). We fit the measured hyperfine resonances between $\mathrm{V}_{\mathrm{B}}^{-}$ $|m_{s}=0\rangle\leftrightarrow|m_{s}=+1\rangle$ in the ODMR spectra as a function of the laser power to extract the nuclear spin polarization fidelity (Fig. 4a). We find that the nuclear spin polarization at both $10$ K and 300300300~{}300K saturate around $30\%$ with similar laser powers ($\sim 10~{}$mW) (Fig. 4b), suggesting that the $\mathrm{V}_{\mathrm{B}}^{-}$ spin lifetime $T_{1}$ and hyperfine interaction strength are not the limiting sources for nuclear polarization.
@@
Figure 4: Dynamic nuclear polarization at low temperature. (a) Near esLAC level ODMR Spectra of the $|m_{s}=0\rangle\leftrightarrow|m_{s}=-1\rangle$ transition at 10 K and 300 K. The spectra both exhibit similar asymmetry toward the left peaks, indicating a polarization of nuclear spins (see Supplemental Material) (b) Polarization level of the three nearest ${}^{15}\mathrm{N}$ nuclear spins near esLAC level as a function of the normalized laser power under room/cryogenic temperature. Inset: schematic representation of the polarization process. The strong spin-conserving optical polarization (green arrow) continuously pumps state from $|m_{s}=-1\rangle$ to $|m_{s}=0\rangle$, and the two hybridisation processes (red arrow and dashed blue arrow) differ in strength, resulting in the polarization of $|m_{I}=\>\uparrow\rangle$ state (see Supplemental Material).
@@
The spin polarization may be limited from the spin non-conserving terms in the hyperfine interaction (see Supplemental Material), which are temperature insensitive. In addition, the observed nuclear spin polarization of $\mathrm{V}_{\mathrm{B}}^{-}$ persists down to 101010~{}10K, in sharp contrast to the case of NV centers where the dynamic nuclear polarization process is substantially suppressed at $T\lesssim 50~{}$K, due to the lack of thermal averaging between NV excited states originating from different orbitals [65, 79]. For $\mathrm{V}_{\mathrm{B}}^{-}$ , the lifting of the double degeneracy of excited states leads to a sufficiently large energy gap between the states, which renders the system insensitive to dynamic Jahn–Teller distortions [78].
@@
Conclusion and Outlook— In summary, our research offers a comprehensive investigation of the temperature-dependent spin-phonon interaction of $\mathrm{V}_{\mathrm{B}}^{-}$ centers in isotopically purified $\mathrm{h}{}^{10}\mathrm{B}{}^{15}\mathrm{N}$. Our results will be of critical importance in designing experiments to make use of $\mathrm{V}_{\mathrm{B}}^{-}$ centers as nanoscale thermometers [47, 57, 55]. While we focus on $\mathrm{V}_{\mathrm{B}}^{-}$, the methodologies developed here can be readily applied to a broad family of spin defects in two-dimensional materials [80, 81, 82, 83]
In addition, our results pave the way for a number of important directions beyond thermometry. First, owing to the 2D nature of the host hBN material, the local spin-phonon interaction of $\mathrm{V}_{\mathrm{B}}^{-}$ is expected to be strongly influenced by interfacial stacking. Therefore, $\mathrm{V}_{\mathrm{B}}^{-}$ centers in a few-layer hBN [61] can potentially offer a brand-new quantum sensing modality — probing interfacial phonon interactions in 2D devices. Second, leveraging the 2D interfaces can provide the opportunity to engineer the local spin-phonon interactions, thereby further enhancing the spin properties of $\mathrm{V}_{\mathrm{B}}^{-}$. We emphasize that these two new directions cannot be pursued by using NV centers in 3D diamond, as the spin-phonon interaction there is confined within the bulk diamond lattice and not expected to be influenced by the diamond surface.
@@
Acknowledgements: We gratefully acknowledge Vincent Jacques, Tongcang Li, Shankar Mukherji, Erdong Song, Xi Wang and Norman Yao for helpful discussions. This work is supported by NSF ExpandQISE 2328837 and the Center for Quantum Leaps of Washington University. B. Huang, Y. Jin and G. Galli acknowledge the support of Air Force Office of Scientific Research (AFOSR) under Contract No. FA9550-22-1-0370, the Next Generation Quantum Science and Engineering (Q-NEXT) hub supported by the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers and resources of the University of Chicago Research Computing Center. Support for hBN crystal growth (E. Janzen and J. H. Edgar) was provided by the Office of Naval Research, award no. N00014-22-1-2582.